The manager of a bulk foods establishment sells a trail mix for $8 per pound and premium cashews for $14 per pound. The manager wishes to make a 420 pound mix that will sell for $9 per pound. How many pounds of each should be used?
Answer by hokies(65)

trail mix = t
cashews = c
t + c = 420
8t + 14c = 9*420
t = 350 pounds
c = 70 pounds